OpenStudy (anonymous):

The graph of (x/2)^2-(y/3)^2=1 is a hyperbola. Which set of equations represents the asymptotes of the hyperbola's graph? A) y=(3/2)x, y=(-3/2)x B) y=(2/3)x, y=(-2/3)x C) y=(1/2)x, y=(-1/2)x D) y=(1/3)x, y=(-1/3)x

OpenStudy (anonymous):

please help(:

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

A hyperbola has an equation (x-h)^2/a^2 + (y-k)^2/b^2 = 1 where h, k are constants and a, b are also constants. So now that you know a, b, h, k you can use the formula for the asymptote: \[y = \pm \frac{ a }{ b } (x - h) + k\]

OpenStudy (anonymous):

So can you identify the variables: a = ? b = ? h = ? k = ?

OpenStudy (anonymous):

ive never done this before so I dont know

OpenStudy (anonymous):

i could take a guess. a=2 b=3 ummm

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

Well the formula you provided is: \[\frac{ x ^{2} }{ 4 } - \frac{ y ^{2} }{ 9 } = 1\] And the formula for a hyperbole is: \[\frac{ (x - h)^{2} }{ a^{2} } + \frac{ (y - k)^{2} }{ b^{2} } = 1\]

OpenStudy (anonymous):

Do you see the similarities there?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

hyperbole?

OpenStudy (anonymous):

Hyperbola* sorry. Also, the formula should say - not +. But see: \[(x - h)^{2} = x^{2}\] \[(y - k)^{2} = y^{2}\] \[a^{2} = 4\] \[b^{2} = 9\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what are a, b, h and k?

OpenStudy (anonymous):

a=2 b=3 h=x/2 k=y/3

OpenStudy (anonymous):

Whoa, h and k are constants. They do not contain x or y in any way. If (x - h)^2 = x^2, what is h?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Yes. Now take your values and plug them in to the asymptote equation: \[y = \pm \frac{ a }{ b } (x - h) + k\]

OpenStudy (anonymous):

did we figure out wht h and k were?

OpenStudy (anonymous):

You found h, now apply the same reasoning to find k.

OpenStudy (anonymous):

so \[\pm2/3(x-0)+k\] ummm idk

OpenStudy (anonymous):

oh k is 0..my bad

OpenStudy (anonymous):

What is the correct answer to the problem?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

That is correct! Good work. Bookmark this thread if you have trouble in future with asymptotes.

OpenStudy (anonymous):

alright. thank you so much